OpenStudy (anonymous):

I need someone to work out this question for me, I already know the answer; I just need to know how to do this.

OpenStudy (anonymous):

The polygons are similar, but not necessarily drawn to scale. Find the value of x. ANSWER: x = 9, y = 27

hartnn (hartnn):

when the polygons are similar, their sides are proportional.

hartnn (hartnn):

so, (x-1)/6 = 16/12 you get x from here.

hartnn (hartnn):

(y+1) /21 = 16/12 you get y from here.

OpenStudy (anonymous):

Ahh, okay. That makes sense! Thank you so, so much.

hartnn (hartnn):

welcome ^_^